OpenStudy (anonymous):

Indicate the equation of the given line in standard form.The line that is the perpendicular bisector of the segment whose endpoints are R(-1, 6) and S(5, 5)

OpenStudy (mathmate):

 Three steps are required: 1. Find the mid-point (M) of the segment using Mid-point (M)of a segment between points \(R(x_1,y_1)\) and \(S(x_2,y_2)\) \(\large M=(\frac{x_1+x_2)}{2},\frac{y_1+y_2}{2})\) 2. Find the slope of segment RS using Slope (m) between two points \(R(x_1,y_1)\) and \(S(x_2,y_2)\) \(\Large m=\frac{y_2-y_1)}{x_2-x_1}\) 3. find the perpendicular bisector, line L through M with slope m. Here's a guide: \(\color{teal}{m} \)= slope \((\color{blue}{x_0},\color{purple}{y_0})\) = given point M Equation of line \(\boxed{\color{red}{L: y-\color{purple}{y_0} = \color{teal}{m} (x-\color{blue}{x_0})}}\) Example: Find equation of the line with slope 2.5 passing through the point (4,3) \(\color{teal}{m} = \color{teal}{2.5}\) \((\color{blue}{x_0},\color{purple}{y_0})=(\color{blue}{4},\color{purple}{3})\) \(L: (y-\color{purple}{3}) = \color{teal}{2.5} (x-\color{blue}{4}) => y-\color{purple}{3}=\color{teal}{2.5}x-10 => y = 2.5x-7\)